OpenStudy (anonymous):

Quadratic equations 4x2 - 16 = 0 2x2 = 16 x2 + 5 = 0 x2 + 4x - 6 = 0 (Solve by completing the square) x2 - 25 = 0 3x2 = -9x 2x2 - 4x + 1 = 0 3x2 + 2x = 7 x2 + 8x + 9 = -7

OpenStudy (dan815):

which one u want to solve?

OpenStudy (dan815):

do u want to know how to complete the square?

OpenStudy (dan815):

x2 + 4x - 6 = 0 x2 + 4x = 6 x2 + 4x +4 = 6 +4 (x+2)(x+2)=10 (x+2)^2=10 x=\[sqrt10 - 2\]

OpenStudy (dan815):

look up youtube video on complete the square if u dont know already

OpenStudy (anonymous):

there all seperate equations

OpenStudy (dan815):

and?